# Actor-Critic for Continuous Action Chunks: A Reinforcement Learning Framework for Long-Horizon Robotic Manipulation with Sparse Reward

### v1

@@ Introduction
For long-horizon tasks, predicting entire sequences of actions—a paradigm known as action chunking—has proven highly successful, finding its most natural application in Imitation Learning (IL). This success is exemplified by recent methods like ACT (Zhao et al. 2023) and Diffusion Policy (Chi et al. 2023), which use powerful generative models to effectively clone complex, continuous action sequences from expert data. Despite their power, these IL methods are constrained by an upper bound, as their performance is limited by the provided expert demonstrations. Due to distributional shift (Ross, Gordon, and Bagnell 2011), the agent is likely to fail when it encounters unseen states that deviate even slightly from the expert trajectory (Fig. 1(a)). The promise of overcoming these limitations, through online interaction and the potential to discover superior policies, provides a motivation to integrate the principles of RL.
However, bringing the chunking paradigm into RL is non-trivial. The complexity of the exploration space increases exponentially as the chunk length grows, making exploration and value estimation intractable for standard methods and frequently leading to exploding $Q$-value estimates (Seo and Abbeel 2025) and subsequent training instability. Recent work CQN-AS (Seo and Abbeel 2025) addressed this by discretizing a library of chunks and learning a $Q$-function over that set, but the discretization inevitably sacrifices precision and flexibility (as shown in Fig. 1(b)). Although other concurrent work (Li, Zhou, and Levine 2025) has attempted to generate continuous action chunks via complex distillation pipelines, these methods often rely on large-scale offline datasets and are computationally expensive, limiting their application in few-shot, real-time control settings. Consequently, a critical gap remains in the field: an efficient RL framework that can directly and data-efficiently leverage the full potential of action chunking in continuous domains.
@@
Imitation Learning from Demonstrations. Imitation Learning (IL) offers a data-driven approach to policy learning by leveraging expert demonstrations, avoiding the exploration challenges of pure RL. The most fundamental method, Behavioral Cloning (BC), treats policy learning as a supervised problem but is known to suffer from distributional shift (Ross, Gordon, and Bagnell 2011). To better address long-horizon tasks, the IL paradigm has evolved towards action chunking, where the policy learns to predict entire sequences of actions. State-of-the-art methods in this domain, such as ACT (Zhao et al. 2023; Lee et al. 2024) and Diffusion Policy (Chi et al. 2023; Ren et al. 2024; Wang et al. 2025), employ powerful generative models (e.g., Transformers (Vaswani et al. 2017) and diffusion models (Ho, Jain, and Abbeel 2020)) to approximate the action distribution, and have been extended to multi-task setups (Bharadhwaj et al. 2024; Liu et al. 2024; Doshi et al. 2024), mobile manipulation (Fu, Zhao, and Finn 2024; Motoda et al. 2025) and humanoid control (Fu et al. 2024). However, their performance is fundamentally capped by the expert data and they cannot improve through online interaction, which motivates integrating the principles of RL.
Reinforcement Learning with Action Chunks. Structured modeling of action sequences is an important approach to improving the efficiency of policy learning and execution (Saanum et al. 2023). However, applying the action chunking paradigm to a RL context is non-trivial, as it dramatically expands the action space for an exploring agent. Discovering an effective policy through exploration in long-horizon, sparse-reward environments is quite difficult, thus existing works generally rely on demonstrations for exploration guidance. ResiP (Ankile et al. 2024) applies RL as a residual component to fine-tune action chunk predicted by BC, but its RL process still performs policy updates at the low-level, single-timestep scale. T-SAC (Tian et al. 2025) learns a critic on action chunks by integrating $n$-step returns while still optimizing a single-timestep actor. CQN-AS (Seo and Abbeel 2025), in contrast, uses the action chunk as the high-level unit for RL exploration. However, its reliance on $Q$-learning forces a discretization of the action space, fundamentally limiting the policy’s precision in continuous control. While more recent work $Q$-Chunking (Li, Zhou, and Levine 2025) overcomes this by generating continuous chunks via a complex distillation pipeline, it introduces computational overhead and relies on large-scale offline datasets. Therefore, developing a framework that can learn continuous action chunks in a direct, stable, and computationally efficient manner remains a challenge. To address this challenge, we introduce AC3.
@@
We model the robotic manipulation task as a semi-Markov Decision Process (SMDP) (Sutton, Precup, and Singh 1999) with continuous chunk space, defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$. Here, $\mathcal{S}$ is a high-dimensional state space, which includes RGB visual states and the robot’s proprioceptive states. $\mathcal{A}$ is a continuous action chunk with fixed length, $P$ is a state transition dynamics, $\gamma\in[0,1)$ is a discount factor, and $r$ is the reward function. For the tasks we consider, the reward $r\in\{0,1\}$ is sparse, providing a positive signal only upon successful task completion. We utilize a small offline dataset of $N$ expert demonstrations $\mathcal{D}_{\text{demo}}=\{\tau_{1},...,\tau_{N}\}$, where each trajectory $\tau$ is a sequence of state-action pairs. Our objective is to learn a policy $\pi_{\theta}(s)$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$.
@@
similar to TD3 (Fujimoto, Hoof, and Meger 2018), the exploration noise $\epsilon$ is sampled from $\text{clip}(\mathcal{N}(0,\sigma),-c,c)$, and we keep two target critics and use their clipped double-Q estimate $\min_{i}Q_{\phi^{\prime}_{i}}$ to curb over-estimation bias and stabilize learning. The critic network still employs a simple MLP+GRU architecture similar to actor network.
@@
for timestep $t=1\text{ to }T$ do   if t $\text{ mod }C t$hen     Compute action chunk with exploration noise $\mathcal{A}_{t}^{\text{exp}}\sim\pi_{\theta}(s_{t})+\epsilon$, $\epsilon\sim\text{clip}(\mathcal{N}(0,\sigma),-c,c)$   end if    Pop action $a_{t}$ from current action chunk, execute $a_{t}$ in the environment, obtain reward $r_{t}$ and new state $s_{t+1}$   $r_{t}\leftarrow r_{\text{int}}(s_{t})$ using Eq. 9 if not terminal $s_{t}$   Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{B}$    Sample $B$ transitions $(s_{t},\mathcal{A}_{t}^{\text{exp}},r_{t:t+n-1},s_{t+n})$ from replay buffer $\mathcal{B}$ as a mini-batch   // Critic Update   Compute target value $y_{t}$ using Eq. 2   $\mathcal{L}_{\phi_{i}}=\frac{1}{B}\sum[(Q_{\phi_{i}}(s_{t},A_{t}^{\text{exp}})-y_{t})^{2}]~~~~\forall i\in\{1,2\}$   $\phi_{i}\leftarrow\phi_{i}-\alpha\nabla_{\phi_{i}}\mathcal{L}_{\phi_{i}}~~~~\forall i\in\{1,2\}$   $\phi^{\prime}_{i}\leftarrow(1-\mu)\phi^{\prime}_{i}-\mu\phi_{i}~~~~\forall i\in\{1,2\}$   // Actor Update    Filter $B_{\text{succ}}$ successful trajectories’ transitions from current mini-batch through successful mask $m_{\text{succ}}$   $\mathcal{L}_{\theta_{BC}}=\frac{1}{B_{\text{succ}}}\sum\left[\|\pi_{\theta}(s_{t})-\mathcal{A}_{t}^{\text{exp}}\|^{2}*m_{\text{succ}}\right]$   $\mathcal{L}_{\theta_{Q}}=-\frac{1}{B_{\text{succ}}}\sum\left[\min_{i=1,2}Q_{\phi_{i}}(s_{t},\pi_{\theta}(s_{t}))*m_{\text{succ}}\right]$   $\mathcal{L}_{\theta}=\lambda_{BC}\mathcal{L}_{\theta_{BC}}+\lambda_{Q}\mathcal{L}_{\theta_{Q}}$   $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta}$ end for
@@
  • RLBench: RLBench (James et al. 2020) is a benchmark focusing on tabletop manipulation. We select 10 tasks, with the same reward setting as BiGym. We use RGB observations with 84×84 resolution from front, wrist, left shoulder, and right shoulder cameras. We maintain the same demonstration collection method as CQN (Seo, Uruç, and James 2024), with 100 demonstrations per task.
@@
Methods
Trainable
Parameters (M)
Inference
Speed (ms)
@@ Conclusion
Input Image
Resolution $\times84$$\times$3
@@
Goal Network
Architecture Figure 9(a)
@@
Critic Network
Architecture Figure 9(b)
@@
Actor Network
Architecture Figure 9(c)
@@
Other Training Setting
Action chunk size ($C$) 16
@@
Special Cases: QC-FQL. Q-Chunking (QC) (Li, Zhou, and Levine 2025) is a concurrent work that proposes a fully chunk-based RL algorithm, i.e., where $n=C$. It employs flow matching (Lipman et al. 2022; Park, Li, and Levine 2025) for behavior cloning pre-training on large-scale offline datasets, followed by online RL using all past experiences. We reproduced the actor-critic version of QC, i.e., QC-FQL, using the PyTorch framework, which can also directly output continuous action chunks. For QC-FQL, we construct the actor_bc_flow and actor_onestep_flow networks using a network architecture similar to the actor in AC3, and construct its value network with an architecture identical to the critic in AC3. We allocated half of the training steps to offline pre-training and the other half to online RL, and the experimental results are presented in Figure 13. QC-FQL-4 indicates a chunk size is 4, and QC-FQL-16 is analogous.

### v2

# arXiv:2508.11143v2[cs.RO] 27 Feb 2026

@@ Introduction
For long-horizon tasks, predicting entire sequences of actions—a paradigm known as action chunking—has proven highly successful, finding its most natural application in Imitation Learning (IL). This success is exemplified by recent methods like ACT (Zhao et al. 2023) and Diffusion Policy (Chi et al. 2023), which use powerful generative models to effectively clone complex, continuous action sequences from expert data. Despite their power, these IL methods are constrained by an upper bound, as their performance is limited by the provided expert demonstrations. Due to distributional shift (Ross et al. 2011), the agent is likely to fail when it encounters unseen states that deviate even slightly from the expert trajectory (Fig. 1(a)). The promise of overcoming these limitations, through online interaction and the potential to discover superior policies, provides a motivation to integrate the principles of RL.
However, bringing the chunking paradigm into RL is non-trivial. The complexity of the exploration space increases exponentially as the chunk length grows, making exploration and value estimation intractable for standard methods and frequently leading to exploding $Q$-value estimates (Seo and Abbeel 2025) and subsequent training instability. Recent work CQN-AS (Seo and Abbeel 2025) addressed this by discretizing a library of chunks and learning a $Q$-function over that set, but the discretization inevitably sacrifices precision and flexibility (as shown in Fig. 1(b)). Although other concurrent work (Li et al. 2025) has attempted to generate continuous action chunks via complex distillation pipelines, these methods often rely on large-scale offline datasets and are computationally expensive, limiting their application in few-shot, real-time control settings. Consequently, a critical gap remains in the field: an efficient RL framework that can directly and data-efficiently leverage the full potential of action chunking in continuous domains.
@@
Imitation Learning from Demonstrations. Imitation Learning (IL) offers a data-driven approach to policy learning by leveraging expert demonstrations, avoiding the exploration challenges of pure RL. The most fundamental method, Behavioral Cloning (BC), treats policy learning as a supervised problem but is known to suffer from distributional shift (Ross et al. 2011). To better address long-horizon tasks, the IL paradigm has evolved towards action chunking, where the policy learns to predict entire sequences of actions. State-of-the-art methods in this domain, such as ACT (Zhao et al. 2023; Lee et al. 2024) and Diffusion Policy (Chi et al. 2023; Ren et al. 2024; Wang et al. 2025), employ powerful generative models (e.g., Transformers (Vaswani et al. 2017) and diffusion models (Ho et al. 2020)) to approximate the action distribution, and have been extended to multi-task setups (Bharadhwaj et al. 2024; Liu et al. 2024; Doshi et al. 2024), mobile manipulation (Fu et al. 2024b; Motoda et al. 2025) and humanoid control (Fu et al. 2024a). However, their performance is fundamentally capped by the expert data and they cannot improve through online interaction, which motivates integrating the principles of RL.
Reinforcement Learning with Action Chunks. Structured modeling of action sequences is an important approach to improving the efficiency of policy learning and execution (Saanum et al. 2023). However, applying the action chunking paradigm to a RL context is non-trivial, as it dramatically expands the action space for an exploring agent. Discovering an effective policy through exploration in long-horizon, sparse-reward environments is quite difficult, thus existing works generally rely on demonstrations for exploration guidance. ResiP (Ankile et al. 2024) applies RL as a residual component to fine-tune action chunk predicted by BC, but its RL process still performs policy updates at the low-level, single-timestep scale. T-SAC (Tian et al. 2025) learns a critic on action chunks by integrating $n$-step returns while still optimizing a single-timestep actor. CQN-AS (Seo and Abbeel 2025), in contrast, uses the action chunk as the high-level unit for RL exploration. However, its reliance on $Q$-learning forces a discretization of the action space, fundamentally limiting the policy’s precision in continuous control. While more recent work $Q$-Chunking (Li et al. 2025) overcomes this by generating continuous chunks via a complex distillation pipeline, it introduces computational overhead and relies on large-scale offline datasets. Therefore, developing a framework that can learn continuous action chunks in a direct, stable, and computationally efficient manner remains a challenge. To address this challenge, we introduce AC3.
@@
We model the robotic manipulation task as a semi-Markov Decision Process (SMDP) (Sutton et al. 1999) with continuous chunk space, defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$. Here, $\mathcal{S}$ is a high-dimensional state space, which includes RGB visual states and the robot’s proprioceptive states. $\mathcal{A}$ is a continuous action chunk with fixed length, $P$ is a state transition dynamics, $\gamma\in[0,1)$ is a discount factor, and $r$ is the reward function. For the tasks we consider, the reward $r\in\{0,1\}$ is sparse, providing a positive signal only upon successful task completion. We utilize a small offline dataset of $N$ expert demonstrations $\mathcal{D}_{\text{demo}}=\{\tau_{1},...,\tau_{N}\}$, where each trajectory $\tau$ is a sequence of state-action pairs. Our objective is to learn a policy $\pi_{\theta}(s)$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$.
@@
similar to TD3 (Fujimoto et al. 2018), the exploration noise $\epsilon$ is sampled from $\text{clip}(\mathcal{N}(0,\sigma),-c,c)$, and we keep two target critics and use their clipped double-Q estimate $\min_{i}Q_{\phi^{\prime}_{i}}$ to curb over-estimation bias and stabilize learning. The critic network still employs a simple MLP+GRU architecture similar to actor network.
@@
for timestep $t=1\text{ to }T$ do   if t $\text{ mod }C t$hen     Compute action chunk with exploration noise $\mathcal{A}_{t}^{\text{exp}}\sim\pi_{\theta}(s_{t})+\epsilon$, $\epsilon\sim\text{clip}(\mathcal{N}(0,\sigma),-c,c)$   end if    Pop action $a_{t}$ from current action chunk, execute $a_{t}$ in the environment, obtain reward $r_{t}$ and new state $s_{t+1}$   $r_{t}\leftarrow r_{\text{int}}(s_{t})$ using Eq. 9 if not terminal $s_{t}$   Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{B}$    Sample $B$ transitions $(s_{t},\mathcal{A}_{t}^{\text{exp}},r_{t:t+n-1},s_{t+n})$ from replay buffer $\mathcal{B}$ as a mini-batch   // Critic Update   Compute target value $y_{t}$ using Eq. 2   $\mathcal{L}_{\phi_{i}}=\frac{1}{B}\sum[(Q_{\phi_{i}}(s_{t},A_{t}^{\text{exp}})-y_{t})^{2}]~~~~\forall i\in\{1,2\}$   $\phi_{i}\leftarrow\phi_{i}-\alpha\nabla_{\phi_{i}}\mathcal{L}_{\phi_{i}}~~~~\forall i\in\{1,2\}$   $\phi^{\prime}_{i}\leftarrow(1-\mu)\phi^{\prime}_{i}+\mu\phi_{i}~~~~\forall i\in\{1,2\}$   // Actor Update    Filter $B_{\text{succ}}$ successful trajectories’ transitions from current mini-batch through successful mask $m_{\text{succ}}$   $\mathcal{L}_{\theta_{BC}}=\frac{1}{B_{\text{succ}}}\sum\left[\|\pi_{\theta}(s_{t})-\mathcal{A}_{t}^{\text{exp}}\|^{2}*m_{\text{succ}}\right]$   $\mathcal{L}_{\theta_{Q}}=-\frac{1}{B_{\text{succ}}}\sum\left[\min_{i=1,2}Q_{\phi_{i}}(s_{t},\pi_{\theta}(s_{t}))*m_{\text{succ}}\right]$   $\mathcal{L}_{\theta}=\lambda_{BC}\mathcal{L}_{\theta_{BC}}+\lambda_{Q}\mathcal{L}_{\theta_{Q}}$   $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta}$ end for
@@
  • RLBench: RLBench (James et al. 2020) is a benchmark focusing on tabletop manipulation. We select 10 tasks, with the same reward setting as BiGym. We use RGB observations with 84×84 resolution from front, wrist, left shoulder, and right shoulder cameras. We maintain the same demonstration collection method as CQN (Seo et al. 2024), with 100 demonstrations per task.
@@
Methods Trainable Parameters (M) Inference Speed (ms)
@@ Conclusion
Input Image Resolution $\times84$$\times$3
@@
Goal Network Architecture Figure 9(a)
@@
Critic Network Architecture Figure 9(b)
@@
Actor Network Architecture Figure 9(c)
@@
Other Training Setting Action chunk size ($C$) 16
@@
Special Cases: QC-FQL. Q-Chunking (QC) (Li et al. 2025) is a concurrent work that proposes a fully chunk-based RL algorithm, i.e., where $n=C$. It employs flow matching (Lipman et al. 2022; Park et al. 2025) for behavior cloning pre-training on large-scale offline datasets, followed by online RL using all past experiences. We reproduced the actor-critic version of QC, i.e., QC-FQL, using the PyTorch framework, which can also directly output continuous action chunks. For QC-FQL, we construct the actor_bc_flow and actor_onestep_flow networks using a network architecture similar to the actor in AC3, and construct its value network with an architecture identical to the critic in AC3. We allocated half of the training steps to offline pre-training and the other half to online RL, and the experimental results are presented in Figure 13. QC-FQL-4 indicates a chunk size is 4, and QC-FQL-16 is analogous.